OpenStudy (anonymous):

Use a graphing utility to graph the two equations in the same viewing window. Use the graphs to determine whether the expressions are equivalent. Verify the results algebraically. PLEASE HELP ME ON THIS. y1 = sin x sec x, y2 = tan x

OpenStudy (jdoe0001):

notice the graph, they do overlap, thus they equal to each other

OpenStudy (jdoe0001):

\(\bf sin(x)sec(x)=tan(x)\\ \quad \\ \quad \\ sin(x)\cdot \cfrac{1}{cos(x)}=tan(x)\implies \cfrac{sin(x)}{cos(x)}=tan(x) \) so... what do you think? are they equal?

OpenStudy (anonymous):

I think they are equivalent.

OpenStudy (anonymous):

Am I right?

OpenStudy (anonymous):

Alright i think its equivalent. thanks!

OpenStudy (jdoe0001):

yeap